The present value today of K payable at the end of 4 years from now
is 81.87 if the compound model is assumed and the interest is credited at a
force of interest of δ.
The present value today of K payable at the end of 6 years from now is 83.53
if the interest is credited at a force of interest of δt = t for t ≥ 0. 100
Calculate δ and the effective annual discount rate which is equivalent to δ.
Answer #1

Present value of K payable at the end of 4 years from now is 81.87

81.87e = K ............equation(1)

Present value of K payable at the end of 6 years from now is 83.53 and δt = t for t ≥ 0

83.53 expdtK 0

83.53 mathbf{e^{[rac{t^{2}}{2}]_{0}^{6}}} = K

83.53 exp[18-0] = K

Substituting value of K from eq(1)

83.53 e[18-0] = 81.87e

e*s 83.53e 81.87

45 = In(-8187 83.53 e18 81.87

δ = 4.505

Annual Discount Rate:

Let d = Annual discount rate

d = 1-e

d = 0.9889